# Semi-adaptive Synergetic Two-way Pseudoinverse Learning System

### v1

@@ Abstract
Deep learning Non-gradient descent learning Pseudoinverse learing Synergetic learning system.
@@ 1 Introduction
Deep learning[12, 18], as a powerful representation learning technology, has achie-ved remarkable accomplishments in various domains, exerting profound influences on human society[20, 34]. Gradient descent algorithm and its variants are a class of commonly used optimization algorithms employed to train deep neural networks by minimizing loss functions. The basic idea is to iteratively adjust parameters in the opposite direction of the gradient of the loss function, gradually approaching the minimum point of the function[25]. These algorithms are widely utilized in the field of deep learning due to their simplicity, ease of implementation and parallelizability[4]. However, gradient descent algorithms face a range of challenges, including low training efficiency, the difficulty in setting hyperparameters, and the possibility of encountering issues like gradient vanishing and gradient explosion. Simultaneously, network architecture design and computational resource constraints represent two significant challenges in deep learning.
To overcome the limitations of gradient descent algorithms, researchers have investigated many non-gradient descent methods. Extreme Learning Machine (ELM)[16, 29] is a model based on a single-hidden layer feedforward neural network (SLFN). Unlike conventional gradient based algorithms such as backpropagation, ELM employs a unique training strategy: the input weights and biases are randomly assigned, and then the output weights are analytically determined using the Moore-Penrose generalized inverse. Hierarchical Extreme Learning Machine (HELM) [27] is an extension of ELM, introducing a hierarchical structure to enhance the model’s capability and performance. Broad learning system (BLS)[3, 5] draws inspiration from the random vector functional link neural network (RVFLNN)[21, 23, 24]. BLS is configured as a flat network, where the initial inputs are embedded into feature nodes. The structure then undergoes broad expansion through enhancement nodes. The key idea of these representative non-gradient descent learning algorithms can also be traced back to the PseudoInverse Learning algorithm (PIL)[7, 10] which was originally proposed for training SLFN. In PIL, the output weights are calculated analytically by calculating an approximate optimal solution for the loss function using the Moore-Penrose generalized inverse. Several variants of the PIL algorithm[30] have been investigated that can use either the random weights or the pseudo-inverse of the input data or its low-rank approximation as input weights.
@@ 2.1 Pseudoinverse Learning based Autoencoder
An autoencoder[15, 22], a commonly used neural network employed in unsupervised learning, comprises an encoder and a decoder. The encoder transforms input data into a low-dimensional representation, while the decoder reconstructs this representation to an output closely resembling the original input. The loss function for an autoencoder can be defined as
@@ 2.2 Synergetic Learning System
A Synergetic Learning System (SLS)[33, 8, 35, 9] is a system that integrates at least two subsystems, which can be agents, models, or human-machine synergies. Synergetic learning between subsystems can be master and servant, cooperative or adversarial, as determined by the learning task.
As shown in Fig. 2, SLS can be built hierarchically, beginning with the assembly of individual neurons into functional units called chunks. These chunks are then integrated into larger subsystems, which, when combined, form the intricate architecture of neural networks. Finally, SLS emerges as a synergetic network comprising multiple interconnected subsystems, each contributing to the overall functionality of the system.
@@ 3.1.1 Forward Learning
During the forward learning process, we utilize PILAE as the foundational block to construct a multi-layer network, namely stacked PILAE. As depicted in Eq. (2), $L_{1}$ regularization can be chosen when feature selection or achieving a sparse solution is desired, with the setting $r=1$. This optimization problem is also termed as LASSO [28]. To solve this optimization problem, various methods can be employed[37], including least angle regression (LAR)[14], iterative shrinkage-thresholding algorithm (ISTA), and alternating direction method of multipliers (ADMM)[2, 11]. This study utilizes the fast iterative shrinkage-thresholding algorithm (FISTA)[1, 36], an efficient variant of ISTA, to tackle the optimization problem. The main distinction between FISTA and ISTA lies in the selection of the starting point of the approximation function in the iteration step. Specifically, the basic iterative steps of FISTA are as follows
@@ 3.1.2 Backward Learning
Forward learning, which primarily relies on unsupervised data reconstruction tasks for feature extraction, inevitably leads to features that are more suited for the reconstruction task, thus potentially hindering performance on downstream learning tasks. Additionally, forward learning fails to fully leverage the information contained in target labels, which is evidently beneficial for extracting features associated with specific learning tasks. This prompts us to propose backward learning. Once the training of the forward learning network is completed, the structure of the network is established. Backward learning employs the same architecture, learning in reverse from the last hidden layer back to the input layer.
@@ 4.1 Data Sets
In our study, we rigorously evaluate the performance of proposed methodologies by employing a wide array of benchmark data sets that are widely recognized within the research community. Specifically, we include classical and contemporary data sets such as MNIST, Fashion-MNIST (F-MNIST), and the NORB data set [19]. Additionally, we utilize a varied selection of data sets from the UCI Machine Learning Repository (http://archive.ics.uci.edu/ml) and the OpenML platform (http://openml.org), which have been commonly used in related works. These data sets include but are not limited to, Abalone, the Advertisement data set, the Gisette data set, the Human Activity Recognition (HAR) data set, the Kin8nm data set, Madelon, Mfeat, Isolet, Occupancy, Yeast, Semeion, Segment, and Spambase.
@@ 4.2 Experimental Design
The performance of our method was compared to five baselines that included HELM [27], PILAE [6], ELM-AE [38], PILLS [31], and BLS [3]. These methods use the non-gradient learning strategy. The primary distinction between them lies in the mapping linked to the input weights. They use least squares or ridge regression for output weight learning. For the MNIST, Fashion-MNIST, and NORB data sets, the structural settings of the baselines take the optimal structure mentioned in the corresponding papers.
@@
LeNet-5 comprises two convolutional layers with 6 and 16 5x5 filters respectively, followed by two max-pooling layers. It has three fully connected layers with 120, 84, and 10 (output) nodes. LeNet-5 was trained using the Adam optimizer, with an initial learning rate of 0.001, batch size of 64, for 30 epochs. ResNet50 is a 50-layer deep residual network constructed from bottleneck residual blocks. It employs 1x1 and 3x3 convolutional filters, with skip connections to facilitate training of such a deep architecture. For ResNet50, we employed the Adam optimizer with an initial learning rate of 0.1 and a batch size of 64. The network was trained for 20 epochs. VGG16 utilizes very small 3x3 convolutional filters, with a substantially increased depth of 16 weight layers. It comprises five max-pooling layers and two fully-connected layers with 4096 nodes each. VGG16 was originally trained using stochastic gradient descent (SGD) with an initial learning rate of 0.001, batch size of 64, for 5 epochs.
@@ 4.3 Result and Analysis
Table 2: Average ranking of our method vs. baselines
HELM PILAE ELM-AE PILLS BLS Ours
3.74 4.68 3.42 3.84 3.89 1.21
Table 2 shows the comparison of our method with the baselines in terms of the average ranking of accuracy in the test set. Obviously, the method we propose is leading the way.

### v2

@@ Abstract
Deep learning Non-gradient descent learning Pseudoinverse learning Synergetic learning system.
@@ 1 Introduction
Deep learning[11], as a powerful representation learning technology, has achie-ved remarkable accomplishments in various domains, exerting profound influences on human society[29]. Gradient descent algorithm and its variants are a class of commonly used optimization algorithms employed to train deep neural networks by minimizing loss functions. The basic idea is to iteratively adjust parameters in the opposite direction of the gradient of the loss function, gradually approaching the minimum point of the function[20]. These algorithms are widely utilized in the field of deep learning due to their simplicity, ease of implementation and parallelizability[3]. However, gradient descent algorithms face a range of challenges, including low training efficiency, the difficulty in setting hyperparameters, and the possibility of encountering issues like gradient vanishing and gradient explosion. Simultaneously, network architecture design and computational resource constraints represent two significant challenges in deep learning.
To overcome the limitations of gradient descent algorithms, researchers have investigated many non-gradient descent methods. Extreme Learning Machine (ELM)[15, 24] is a model based on a single-hidden layer feedforward neural network (SLFN). Unlike conventional gradient based algorithms such as backpropagation, ELM employs a unique training strategy: the input weights and biases are randomly assigned, and then the output weights are analytically determined using the Moore-Penrose generalized inverse. Hierarchical Extreme Learning Machine (HELM) [22] is an extension of ELM, introducing a hierarchical structure to enhance the model’s capability and performance. Broad learning system (BLS)[2, 4] draws inspiration from the random vector functional link neural network (RVFLNN)[18, 19]. BLS is configured as a flat network, where the initial inputs are embedded into feature nodes. The structure then undergoes broad expansion through enhancement nodes. The key idea of these representative non-gradient descent learning algorithms can also be traced back to the PseudoInverse Learning algorithm (PIL)[6, 9] which was originally proposed for training SLFN. In PIL, the output weights are calculated analytically by calculating an approximate optimal solution for the loss function using the Moore-Penrose generalized inverse. Several variants of the PIL algorithm[25] have been investigated that can use either the random weights or the pseudoinverse of the input data or its low-rank approximation as input weights.
@@ 2.1 Pseudoinverse Learning based Autoencoder
An autoencoder[14], a commonly used neural network employed in unsupervised learning, comprises an encoder and a decoder. The encoder transforms input data into a low-dimensional representation, while the decoder reconstructs this representation to an output closely resembling the original input. The loss function for an autoencoder can be defined as
@@ 2.2 Synergetic Learning System
A Synergetic Learning System (SLS)[7, 8, 30, 28] is a system that integrates at least two subsystems. Subsystems can be agents, models, or human-machine synergies. Synergetic learning between subsystems can be master and servant, cooperative or adversarial, as determined by the learning task.
As shown in Fig. 2, SLS can be built hierarchically, beginning with the assembly of individual neurons into functional units called chunks. These chunks are then integrated into larger subsystems, which combine to form the intricate architecture of neural networks. Finally, SLS emerges as a synergetic network comprising multiple interconnected subsystems, each contributing to the overall functionality of the system.
@@ 3.1.1 Forward Learning
During the forward learning process, we utilize PILAE as the foundational block to construct a multi-layer network, namely stacked PILAE. As depicted in Eq. (2), $L_{1}$ regularization can be chosen when feature selection or achieving a sparse solution is desired, with the setting $r=1$. This optimization problem is also termed as LASSO [23]. To solve this optimization problem, various methods can be employed[32], including least angle regression (LAR)[13], iterative shrinkage-thresholding algorithm (ISTA), and alternating direction method of multipliers (ADMM)[10]. This study utilizes the fast iterative shrinkage-thresholding algorithm (FISTA)[1, 31], an efficient variant of ISTA, to tackle the optimization problem. The main distinction between FISTA and ISTA lies in the selection of the starting point of the approximation function in the iteration step. Specifically, the basic iterative steps of FISTA are as follows
@@ 3.1.2 Backward Learning
Forward learning, which primarily relies on unsupervised data reconstruction tasks for feature extraction, inevitably leads to features that are more suited for reconstruction tasks, thus potentially hindering performance on downstream learning tasks. Additionally, forward learning fails to fully leverage the information contained in target labels which is evidently beneficial for extracting features associated with specific learning tasks. This prompts us to propose backward learning. Once the training of the forward learning network is completed, the structure of the network is established. Backward learning employs the same architecture, learning in reverse from the last hidden layer back to the input layer.
@@ 4.1 Data Sets
In our study, we rigorously evaluate the performance of proposed methodologies by employing a wide array of benchmark data sets that are widely recognized within the research community. Specifically, we include classical and contemporary data sets such as MNIST, Fashion-MNIST (F-MNIST), and the NORB data set [17]. Additionally, we utilize a varied selection of data sets from the UCI Machine Learning Repository (http://archive.ics.uci.edu/ml) and the OpenML platform (http://openml.org), which have been commonly used in related works. These data sets include but are not limited to, Abalone, Advertisement, Gisette, the Human Activity Recognition (HAR) data set, the Kin8nm data set, Madelon, Mfeat, Isolet, Occupancy, Yeast, Semeion and Spambase.
@@ 4.2 Experimental Design
The performance of our method was compared to five baselines that included HELM [22], PILAE [5], ELM-AE [33], PILLS [26], and BLS [2]. These methods use the non-gradient learning strategy. The primary distinction between them lies in the mapping linked to the input weights. They use least squares or ridge regression for output weight learning. For the MNIST, F-MNIST, and NORB data sets, the structural settings of the baselines take the optimal structure mentioned in the corresponding papers.
@@
LeNet-5 comprises two convolutional layers with 6 and 16 $5\times 5$ filters respectively, followed by two max-pooling layers. It has three fully connected layers with 120, 84, and 10 (output) nodes. LeNet-5 was trained using the Adam optimizer, with an initial learning rate of 0.001, batch size of 64, for 30 epochs. ResNet50 is a 50-layer deep residual network constructed from bottleneck residual blocks. It employs $1\times 1$ and $3\times 3$ convolutional filters, with skip connections to facilitate training of such a deep architecture. For ResNet50, we employed the Adam optimizer with an initial learning rate of 0.1 and a batch size of 64. The network was trained for 20 epochs. VGG16 utilizes very small $3\times 3$ convolutional filters, with a substantially increased depth of 16 weight layers. It comprises five max-pooling layers and two fully-connected layers with 4096 nodes each. VGG16 was originally trained using stochastic gradient descent (SGD) with an initial learning rate of 0.001, batch size of 64, for 5 epochs.
@@ 4.3 Result and Analysis
Figure 5: The accuracy ranking box plots of our method and baselines
Fig. 5 shows the comparison of our method with the baselines in terms of ranking of accuracy in the test set. The red vertical line indicates the median and the green triangle indicates the mean. Obviously, the method we propose is leading the way.